# Beyond Statistical Estimation: Differentially Private Individual Computation via Shuffling

### v2

@@ Abstract
In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge. Technologies like Differential Privacy (DP) have been pivotal in addressing these concerns. The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling. These benefits have led to significant interest in the shuffle model. However, the computation tasks in the shuffle model are limited to statistical estimation, making the shuffle model inapplicable to real-world scenarios in which each user requires a personalized output. This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations. PIC enables personalized outputs while preserving privacy, and enjoys privacy amplification through shuffling. We propose a concrete protocol that realizes PIC. By using one-time public keys, our protocol enables users to receive their outputs without compromising anonymity, which is essential for privacy amplification. Additionally, we present an optimal randomizer, the Minkowski Response, designed for the PIC model to enhance utility. We formally prove the security and privacy properties of the PIC protocol. Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
Index Terms:
Differential Privacy, Private Individual Computation, Shuffle Model, Privacy-Utility Trade-offs
I Introduction
@@
Recently, within the realm of DP research, the shuffle model [13, 32] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to a computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [34, 35]. Consequently, to achieve a predefined global privacy goal, less noise must be added when users sanitize their data locally. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [8, 40, 41, 41, 42].
@@
  • III.
    Incentive mechanisms: In federated learning [92] or data crowdsensing [74], incentives play a vital role to encourage well-behaved participation. The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [71]).
At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification. However, this is not necessarily the case. Our observation is that many individual computing tasks are equivariant to shuffling, meaning that the permutation applied to the inputs does not affect the computation. Therefore, shuffling does not prevent the server from producing personalized answers for each client – the server does not need to know which answer belongs to whom. Nevertheless, there is a challenge: how to return the output to the correct user without compromising anonymity. One straightforward approach is to have the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [44]). However, this method is costly due to the need to store communication states and may be vulnerable to de-anonymization attacks on anonymous channels [64, 66]. Additionally, current shuffler implementations, e.g. the one described in the seminal work [13], do not support such duplex connections. Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which is intractable in the duplex shuffle channel. Addressing this issue is the first technical challenge we need to overcome.
The second challenge we face is designing optimal randomizers for individual computation within the shuffle model. Often, the randomizer in a shuffle model protocol should be tailored to specific tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as demonstrated in histogram estimation [39, 34] and one-dimensional summation estimation [7]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population. This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [88]), as well as prevalent randomizers (such as adding Laplace [31] or Staircase [38] noises), less effective for the new setting. Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
@@
Another main contribution is the development of an asymptotically optimal randomizer specifically designed for the PIC model. This randomizer is based on an LDP mechanism we call Minkowski Response. The primary goal of this design is to enhance utility, measured by the single-report error, which is the expected squared error between a user’s true data value and its sanitized version. To achieve this, a Minkowski distance $r$ is determined based on the privacy budget. The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$. The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere in the entire output domain with a relatively low probability.
@@
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC protocol. Section VI provides optimal user-side randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
II Related Works
@@
II-A Secure Multiparty Computation
@@
II-B Curator and Local DP Methods
@@ II-C Shuffle Model of DP
Overall, existing works in the shuffle model primarily focus on statistical queries. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms).
II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [45]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [63] against polynomial-time adversaries. Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table IX).
III Preliminaries
In this section, we provide a concise introduction to the preliminaries. A list of notations used throughout the paper can be found in Table I.
TABLE I: List of notations.
@@ III Preliminaries
$d$ the dimension of user’s data
$\mathbb{X}$ the domain of input data
$\mathbb{Y}$ the domain of a sanitized message
@@
$sk$, $pk$ the secret key and the public key, respectively
@@
III-A Privacy Definitions
Definition III.1 (Hockey-stick divergence [72])
@@ III-A Privacy Definitions
Differential privacy imposes divergence constraints on output probability distributions with respect to changes in the input. In the curator model of differential privacy, a trusted party collects raw data $x_{i}\in\mathbb{X}$ from all users to form a dataset $T=\{x_{1},\ldots,x_{n}\}$ and applies a randomization algorithm $\mathcal{R}$ to release query results $\mathcal{R}(T)$. For two datasets $T$ and $T^{\prime}$ of the same size and differing in only one element, they are referred to as neighboring datasets. Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$). The formal definition of curator differential privacy is as follows:
Definition III.2 ($(\epsilon,\delta)$-DP [31])
@@
In the local model of differential privacy, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition III.3).
Definition III.3 (local $\epsilon$-DP [54])
A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
Definition III.4 (Data processing inequality)
$D(g(P)\|g(Q))\leq D(P\|Q).$
III-B The Classical Shuffle Model
@@
In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [34, 35] have derived that $n$ shuffled $\epsilon_{l}$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP. We denote the amplified privacy level as:
@@
the tight value of which can also be numerically computed [57, 87].
III-C Public Key Encryption
@@
IV Problem Settings
We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the differential privacy sense) for a wide range of data-driven computational tasks. (2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large datasets and user bases. (3) The privacy amplification effect results in significantly better utility compared to direct sanitization and computation using Local Differential Privacy (LDP). In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
IV-A Motivating Applications
@@
Spatial crowdsourcing. A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server (the service platform). It proceeds through the following major steps:
@@
    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
@@
    Querying: Each participating user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which includes information like location $l_{i}$ and preferences $v_{i}$.
@@
    Generating recommendation: The server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria, such as proximity and user preferences.
@@
    Retrieving: Finally, each user retrieves the recommendation results.
@@
    Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [71]).
@@ IV-B The Ideal Functionality
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of users and one server $S$. It sanitizes the input from each user, shuffles the inputs randomly, and then applies a function to compute the output for each user. In the end, each user receives their individual function output, while the server receives a list of the shuffled sanitized user data and a list of the function outputs. It captures the functional requirements of individual computation in the real world: a server performs a computing task using the joint inputs from a set of users (which are sanitized and shuffled), and each user receives an individualized output. It also captures the security requirements: each party receives precisely the specified output, and nothing more.
V A Concrete Protocol
V-A The Protocol
We now present a concrete protocol for PIC. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
@@ V-A The Protocol
Remark 3 To eventually accomplish the spatial crowdsourcing (e.g., taxi-hailing services) or location-based social system tasks, the computation result $y_{i,\pi_{i}(j)}$ will contain the matched users of the user $\pi_{i}(j)\in G_{i}$. That is, $y_{i,\pi_{i}(j)}$ will encapsulate a list of public keys and noisy location information about the matched users of the user $\pi_{i}(j)\in G_{i}$. After receiving the individualized computation results in the final step, each user $\pi_{i}(j)$ can then securely contact the matched users using the public keys in $y_{i,\pi_{i}(j)}$ (possibly with the help of a public communication channel, such as a public bulletin).
V-B Security Analysis
@@
Theorem V.1 (Security)
The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
The proof can be found in Appendix -A. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
@@
Theorem V.2 (Differential Privacy)
@@
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 8(e^{\epsilon}+1)\log(2/\delta)$ we have:
$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest user locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [34, 35]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -B.
Remark 4 After decryption, the server obtains $pk_{i,j}||x^{\prime}_{i,j}$, where $pk_{i,j}$ is a public key not sanitized by the local randomizer. Despite the presence of the public key, $pk_{i,j}||x^{\prime}_{i,j}$ and $x^{\prime}_{i,j}$ are equivalent in terms of privacy amplification. This is because (1) the public key is random and generated independently of $x_{i,j}$, so prepending it to $x^{\prime}_{i,j}$ does not affect the local DP guarantee—it is the same as $x^{\prime}_{i,j}$ itself; and (2) all public keys follow an identical probability distribution across all users in the group $G_{i}$, ensuring that the privacy amplification effect via shuffling is not degraded.
V-C Discussion on Post-computation Communication

### v4

@@ Abstract
In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge. Technologies like differential privacy have been pivotal in addressing these concerns. The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling. These benefits have led to significant interest in the shuffle model. However, the computation tasks in the shuffle model are limited to statistical estimation, making it inapplicable to real-world scenarios in which each user requires a personalized output. This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations. PIC enables personalized outputs while preserving privacy, and enjoys privacy amplification through shuffling. We propose a concrete protocol that realizes PIC. By using one-time public keys, our protocol enables users to receive their outputs without compromising anonymity, which is essential for privacy amplification. Additionally, we present an optimal randomizer, the Minkowski Response, designed for the PIC model to enhance utility. We formally prove the security and privacy properties of the PIC protocol. Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
1 Introduction
@@
[…]
@@
  • Incentive mechanisms: In federated learning [97] or crowd sensing [78], incentives play a vital role to encourage well-behaved participation. The amount of rewarding incentive must be computed for individuals based on their contribution (e.g., via Shapley values [75]).
At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification. However, this is not necessarily the case. Our observation is that many individual computing tasks are equivariant to shuffling, meaning that the permutation applied to the inputs does not affect the computation. Therefore, shuffling does not prevent the server from producing personalized answers for each client – the server does not need to know which answer belongs to whom. Nevertheless, there is a challenge: how to return the output to the correct user without compromising anonymity. One straightforward approach is to have the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [45]). However, this method is costly due to the need to store communication states and may be vulnerable to de-anonymization attacks on anonymous channels [68, 70]. Additionally, current shuffler implementations, e.g. the one described in the seminal work [13], do not support such duplex connections. Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which can be hard to implement in the duplex shuffle channel. Addressing this issue is the first technical challenge we need to overcome.
The second challenge we face is designing optimal randomizers for individual computation within the shuffle model. Often, the randomizer in a shuffle model protocol should be tailored to specific tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as demonstrated in histogram estimation [40, 34] and one-dimensional summation estimation [7]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population. This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, or budget splitting, as reviewed in [93]), as well as prevalent randomizers (such as adding Laplace noise [31]), less effective for the new setting. Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
@@
Another main contribution is the development of an asymptotically optimal randomizer specifically designed for the PIC model. This randomizer is based on an LDP mechanism we call Minkowski Response. The primary goal of this design is to enhance utility, measured by the single-report error, which is the expected squared error between a user’s true data value and its sanitized version. To achieve this, a Minkowski distance $r$ is determined based on the privacy budget. The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$. The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere with a relatively low probability.
@@
Organization. The remainder of this paper is organized as follows. Section 2 reviews related works. Section 3 provides preliminary knowledge about privacy definitions and security primitives. Section 4 formalizes the problem setting. Section 5 presents the PIC protocol. Section 6 provides optimal randomizers. Section 7 evaluates the performances of our proposals. Section 8 discusses more merits and future directions of the PIC. Finally, Section 9 concludes the paper.
2 Related Works
@@
2.1 Secure Multiparty Computation
@@
2.2 Curator and Local DP Methods
@@ 2.3 Shuffle Model of DP
Overall, existing works in the shuffle model primarily focus on statistical queries. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, social systems, and incentive mechanisms).
2.4 Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [46]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [67] against polynomial-time adversaries. Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table 9 of the full version [91]).
3 Preliminaries
In this section, we provide a concise introduction to the preliminaries. A list of notations can be found in Table 1.
Table 1: List of notations.
@@ 3 Preliminaries
$sk$, $pk$ the secret key and public key, respectively
@@
3.1 Privacy Definitions
Definition 3.1 (Hockey-stick divergence [76]).
@@ 3.1 Privacy Definitions
Differential privacy imposes divergence constraints on output probability distributions with respect to changes in the input. In the curator model of differential privacy, a trusted party collects raw data $x_{i}\in\mathbb{X}$ from all users to form a dataset $T=\{x_{1},\ldots,x_{n}\}$ and applies a randomization algorithm $\mathcal{R}$ to release query results $\mathcal{R}(T)$. For two datasets $T$ and $T^{\prime}$ of the same size and differing in only one element, they are referred to as neighboring datasets. Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$).
Definition 3.2 ($(\epsilon,\delta)$-DP [31]).
@@
In the local DP model, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition 3.3).
Definition 3.3 (local $\epsilon$-DP [57]).
A mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
3.2 The Classical Shuffle Model
@@
In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [34, 35] have derived that $n$ shuffled $\epsilon$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP. We denote the amplified privacy level as:
@@
the tight value of which can also be numerically computed [60, 92]. Our experiments will use numerical bounds in [60].
3.3 Public Key Encryption
@@
4 Problem Settings
We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the DP sense) for a wide range of computation tasks. (2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large user bases. (3) The privacy amplification effect results in significantly better utility compared to local DP. In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
4.1 Motivating Applications
@@
Spatial crowdsourcing. A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server. It proceeds through the following major steps:
@@
    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
@@
    Querying: each user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which might include location $l_{i}$ and preferences $v_{i}$.
@@
    Generating recommendation: the server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria (e.g., proximity and preferences).
@@
    Retrieving: each user retrieves the recommendations.
@@
    Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$.
@@ 4.2 The Ideal Functionality
5 A Concrete Protocol
5.1 The Protocol
[…]
We now present a concrete protocol for PIC in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. The $\mathcal{F}$-hybrid model in MPC is a conceptual framework that simplifies the design and analysis of secure protocols. In many MPC protocols, parties rely on existing sub-protocols that have already been proven secure for specific tasks. To help protocol designers focus on other aspects of the protocol, these sub-protocols can be abstracted as ideal functionalities, denoted by $\mathcal{F}$, which are assumed to be computed securely by a trusted third party. This results in a hybrid protocol, combining concrete cryptographic operations from the real world with calls to ideal functionalities, thus avoiding the complexity of specifying and analyzing the sub-protocols in detail. In our case, we define the PIC protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, where $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ replaces the secure shuffle sub-protocol.We now present a concrete protocol for PIC. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
@@ 5.1 The Protocol
[…]
5.2 Security Analysis
@@
Theorem 5.1 (Security).
The Private Individual Computation (PIC) protocol in §5 securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution. The detailed proof can be found in Appendix G of the full version [91].
@@
Theorem 5.2 (Differential Privacy).
@@
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 16e^{\epsilon}\log(2/\delta)$ we have:
$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{64e^{\epsilon}\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{8e^{\epsilon}}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
[…]
5.3 Discussion on Post-computation Communication